Determine the rate law and the value of k for the following reaction using the data provided.

2 N2O5(g) → 4 NO2(g) + O2(g) [ N2O5]i (M) Initial Rate (M-1s-1)
0.033​ 1.84 × 10-4
0.066 7.37 × 10-4
0.099 1.66× 10-3
  • A.

    Rate = 5.6 × 10-1 M-1/2s-1[N2O5]3/2

  • B.

    Rate = 1.7 × 10-1 M-1s-1[N2O5]2

  • C.

    Rate = 1.7 × 10-1M1/2s-1[N2O5]1/2

  • D.

    Rate = 6.0 × 10-1 M-2s-1[N2O5]3

  • E.

    Rate = 5.2 × 10⁻1 s-1[N2O5]
Answer #1

Say the rate equation is,

rate = k. N205)

where, m is the order of the reaction, k is the rate constant, and [ ] stands for concentration.

from the three data,

1.84\times 10^{-4}\, M^{-1}.s^{-1}=k.(0.033\, M)^m..........(i)

7.37\times 10^{-4}\, M^{-1}.s^{-1}=k.(0.066\, M)^m..........(ii)

1.66\times 10^{-3}\, M^{-1}.s^{-1}=k.(0.099\, M)^m..........(iii)

(ii) / (i) gives,

\frac{7.37\times 10^{-4}}{1.84\times 10^{-4}}=\frac{k.0.066^m}{k.0.033^m}

\Rightarrow 4=2^m

\Rightarrow 2^2=2^{m}

\therefore m=2

Thus, rate = k [N2O5]2

Now, applying this in equation (iii) we get,

1.66\times 10^{-3}\, M^{-1}.s^{-1}=k.(0.099\, M)^{2}

k=\frac{1.66\times 10^{-3}\, M^{-1}.s^{-1}}{0.099^2\, M^2}=1.7\times 10^{-1}M^{-3}.s^{-1}

Thus,

rate=1.7\times 10^{-1}\, M^{-3}.s^{-1}\: [N_2O_5\, M]^2=1.7\times 10^{-1}\, M^{-1}.s^{-1}\: [N_2O_5]^2

Thus, asnwer is B.